OpenStudy (anonymous):

Solve the equation. 0.25 r – 0.125 + 0.5 r = 0.5 + r

OpenStudy (anonymous):

You want all 'r's on one side

OpenStudy (anonymous):

0.25r-0.125+0.5r=0.5+r 0.25r+0.5r-r=0.5+0.125 0.75-r=0.625 -r=-0.125 r=0.125